determine whether the following equations represent a parabola, circle, ellipse or hyperbola
4x^2 + 6y^2 - 4x - 9y + 12 = 0
3x^2 + 2x - 5y + 12 = 0
4x^2 + 4y^2 -4x + 9y + 1 = 0
9x^2 + 12 y^2 + 4x + 4y + 4 = 0
5x^2 - 2y^2 + 2x - 6y - 6 = 0
How do I determine this?

Answer by Alan3354(21615)

determine whether the following equations represent a parabola, circle, ellipse or hyperbola
4x^2 + 6y^2 - 4x - 9y + 12 = 0
A theoretical ellipse. It has x^2 and y^2 terms added with differing coefficients. However, the axes are less than zero. If the last term is -12, then it's an ellipse.
--------------------------
3x^2 + 2x - 5y + 12 = 0
A parabola. Only x is 2nd degree.
------------------
4x^2 + 4y^2 -4x + 9y + 1 = 0
A circle. x^2 and y^2 with coefficient 4.
-------------------
9x^2 + 12 y^2 + 4x + 4y + 4 = 0
A theoretical ellipse. It has x^2 and y^2 terms added with differing coefficients. However, the axes are less than zero. If the last term is -4, then it's an ellipse.
-----------------------------------
5x^2 - 2y^2 + 2x - 6y - 6 = 0
A hyperbola. The x^2 and y^2 terms have different signs.
------------------------
How do I determine this?